OpenStudy (anonymous):

write the equation of a line that goes through points (1,1) and (2,4)

OpenStudy (anonymous):

First find the slope using the equation y2 - y1 / x2 - x1, which would be 4 - 1 / 2 - 1, which is 3/1, or 3, so m = 3. Now, use the equation y = mx + b, m being the slope, x and y being a set of coordinates. You can use either set of coordinates to solve for b, which is your y-intercept. So, you will have 4 = 2(3) + b, or 4 = 6 + b Use inverse operations to solve for b. You will get b = 4 - 6, or -2. Now, substitute the m and b into the equation to get your answer, which will be y = 3x - 2.